# Nonlinear Computation with Linear Optics via Source-Position Encoding

### v1

@@ 2.1 Source positional encoding
$c_{m}(\omega)\omega_{m}^{2}=\omega^{2}c_{m}(\omega)-j\omega\braket{\mathcal{E}_{m}(\mathbf{r})}{\mathcal{J}(\mathbf{r},\omega)}_{\epsilon}.$ (6)
@@
Concretely, consider an eigenfunction $\mathcal{E}_{m}$, and suppose it is non-degenerate (the multiplicity of $\omega_{m}$ is $1$), with source frequency $\omega_{s}$ close to its associated eigenfrequency $\omega_{m}$. Then the field is dominated by this particular mode:
@@ 2.2.4 Decoding
MNIST [^2] See (cref Lecun). 8 46.8% 91.7% 82.6%
MNIST 24 73.0% 94.4% 93.1%
MNIST 64 83.3% 95.2% 95.6%
Table 1: Experimental results for benchmark datasets. The linear model is fit numerically using the same quantized inputs as used in the optical neural network. The optical networks optimize the topology in a design space of $500\times 500$, the results cited here are from the neural field parameterization. Additional training and experiment details can be found in Section A.4
@@ 3 Experiments
Our experimental setup is the inverse design of a two dimensional optical classifier using a regular array of cylindrical (i.e., dipole) sources for our positional encoding (see Figure 2). Like prior work (Khoram et al., 2019; Wanjura & Marquardt, 2024), we consider multi-class classification over the MNIST dataset, which is comprised of $28\times 28$ grayscale images at $8$-bit precision, each of which belongs to one of 10 distinct classes.
@@ 3.1 Encoding
We first linearly project each input $\mathbf{x}\in\mathbb{R}^{784}$ to a $p$-vector $\tilde{\mathbf{x}}\in\mathbb{R}^{p}$, with values of $p=4,8,16$ evaluated in our experiments. The linear projection is derived from principal components analysis (PCA). Then, for each of $p$ dimensions comprising $\tilde{\mathbf{x}}$, we uniformly quantize the entry into $q$ bins, reducing the precision of the entry to $\log_{2}q$ bits. In our experiments, we use $q=4,8,16$. These configurations of $(p,q)$ correspond with each image encoded with $8,24$, and $64$ bits. Physically, this encoding corresponds with a $p\times q$ array of identical cylindrical sources laid out within the design area. For any given $\tilde{\mathbf{x}}$, the $p$th row of the source array has precisely one source activated, corresponding with the bin that index is quantized to.
Figure 4: Validation accuracy (as a percentage) against training iteration for the 64 bits/image quantization configuration, shown enveloped by $\pm$ 2 standard deviations. The MLP and linear model best validation accuracies are illustrated as dashed horizontal lines in orange and green respectively.
@@ 3.2 Ablation
Our optical neural networks significantly improve classification performance compared to a linear model (95.2% for the optical network compared to 83.3% for the linear model, in the highest-precision regime), and achieve performance commensurate to the deep artificial neural network.
The artificial neural network is out-performed by the optical network at low precision (see Figure 4), but it is important to note that we use a basic MLP; fitting an optimized, state of the art convolutional neural network (CNN) would almost certainly outperform our optical networks. We also show examples of how the topology evolves over the course of training/optimization in Figure 5. Additional details on experiments and optimization can be found in Section A.4.
@@ A.2 Electromagnetics
$\mathbf{D}:\mathbb{R}^{3}\mapsto\mathbb{C}^{3}$, then $\mathcal{D}(\mathbf{x},t)=\text{Re}(e^{j\omega t}\mathbf{D}(\mathbf{x}))$. Units [C/$\text{m}^{2}$].
@@
$\mathbf{B}:\mathbb{R}^{3}\mapsto\mathbb{C}^{3}$, then $\mathcal{B}(\mathbf{x},t)=\text{Re}(e^{j\omega t}\mathbf{B}(\mathbf{x}))$. Units [W/$\text{m}^{2}$].
@@
$\mathbf{B}:\mathbb{R}^{3}\mapsto\mathbb{C}^{3}$, then $\mathcal{B}(\mathbf{x},t)=\text{Re}(e^{j\omega t}\mathbf{B}(\mathbf{x}))$. Units [W/$\text{m}^{2}$].
@@
$\mathbf{Q}:\mathbb{R}^{3}\mapsto\mathbb{C}$, then $\mathcal{Q}(\mathbf{x},t)=\text{Re}(e^{j\omega t}\mathbf{Q}(\mathbf{x}))$. Units [C/$\text{m}^{3}$].
@@ A.4.1 Does Codesign Matter?
During optimization, we track the root mean square (RMS) of the gradient of the objective with respect to (a) the topology parameters and (b) the decoder parameters, in an attempt to understand the relative contributions to the designs. In general (but with some exceptions), we find that the topology derivatives tend to be between 1 and 3 orders of magnitude larger in RMS than the decoder derivatives; to first order then, the interpretation is that the classification accuracy of the system is around 10-1000x more sensitive to changes in the topology than that of the decoder parameters.
However, it should not be inferred from this quantitative result that the decoder is not critical to the overall system performance. In fact, optimizing without using the decoder at all degrades the system performance significantly. Our interpretation of the relatively smaller decoder gradient magnitude follows the apparent role of the decoder as a ‘permuting’ map to account for the arbitrary assignment of class labels to output measurement locations. Then, after initialization, optimizing the decoder amounts to merely breaking symmetry in its weights to achieve a better class label to output region assignment. From that point, the absolute magnitude of any given decoder weight has minimal effect, since only their relative values are responsible for the label/output region re-mapping.
Optimizing only the decoder but not the topology results in significantly worse performance (around 10% reduction in classification accuracy). But, the fact that the decoder-only performance is significantly better than random indicates that even randomly initialized topologies can be adapted for use as effective classifiers with an appropriately optimized linear decoder; similar to the use of multiple-scattering devices as reservoir networks.

### v2

@@ 2.1 Source positional encoding
$c_{m}(\omega)\omega_{m}^{2}=\omega^{2}c_{m}(\omega)-j\omega\braket{\mathcal{E}_{m}(\mathbf{r})|\mathcal{J}(\mathbf{r},\omega)}_{\epsilon}.$ (6)
@@
Concretely, consider an eigenfunction $\mathcal{E}_{m}$, and suppose it is non-degenerate (the multiplicity of $\omega_{m}$ is 111), with source frequency $\omega_{s}$ close to its associated eigenfrequency $\omega_{m}$. Then the field is dominated by this particular mode:
@@ 2.2.4 Decoding
MNIST [^2] See (cref Lecun). 8 48.6% 56.0 % (+15.2%) 56.6 %
MNIST 24 75.8 % 83.1% (+9.6%) 89.1%
MNIST 64 84.8% 86.7% (+2.2%) 96.9%
FashionMNIST 8 47.1% 53.1% (+12.7%) 54.7%
FashionMNIST 24 64.1% 69.6% (+8.6%) 72.4%
FashionMNIST 64 75.6% 79.1% (+4.6%) 82.1%
Table 1: Experimental results for benchmark datasets. The linear model and MLP are fit numerically using the same quantized data as in the optical neural network. In bold, we provide the relative percentage improvement (over linear) for the optical network configurations. The optical networks optimize the topology in a design space of $\times 500$ 500, the results cited here are from the neural field parameterization. Additional training and experiment details can be found in Section A.4
@@ 3 Experiments
Our experimental setup is the inverse design of a two dimensional optical classifier using a regular array of cylindrical (i.e., dipole) sources for our positional encoding (see Figure 2). Like prior work (Khoram et al., 2019; Wanjura & Marquardt, 2024), we consider multi-class classification over the MNIST dataset, which is comprised of $\times 28$ 28 grayscale images at 888-bit precision, each of which belongs to one of 10 distinct classes.
@@ 3.1 Encoding
[…]
@@ 3.2 Ablation
Our optical neural networks significantly improve classification performance compared to a linear model, and achieve performance generally commensurate to the deep artificial neural network. The numeric deep network outperforms the optical network more at higher precision.
We also show examples of how the topology evolves over the course of training/optimization in Figure 4. Additional details on experiments and optimization can be found in Section A.4.
@@ A.2 Electromagnetics
[…]
@@
[…]
@@
[…]
@@
[…]
@@ A.4.1 Does Codesign Matter?
During optimization, we track the root mean square (RMS) of the gradient of the objective with respect to (a) the topology parameters and (b) the decoder parameters, in an attempt to understand the relative contributions to the designs. In general (but with some exceptions), we find that the using the implicit neural field methods, topology derivatives tend to be between 1 and 3 orders of magnitude larger in RMS than the decoder derivatives; to first order then, the interpretation is that the classification accuracy of the system is around 10-1000x more sensitive to changes in the topology than that of the decoder parameters. That said, we have also found that for many initializations of the topology, the system can be trained in something like a ‘reservoir-mode’, updating only the decoder parameters. It is challenging to state precisely what properties must obtain for a set of topologies to be amenable to this, but this could be explored in future work. At least conceptually, following our modal analysis from Section 2.1, it is relevant to have a methodology for local optimization of the topology.
Optimizing without using the decoder at all degrades the system performance significantly. Our interpretation of the relatively smaller decoder gradient magnitude follows our hypothesis that the decoder acts as a ‘permuting’ map to account for the arbitrary assignment of class labels to output measurement locations. Then, after initialization, optimizing the decoder amounts to merely breaking symmetry in its weights to achieve a better class label to output region assignment. From that point, the absolute magnitude of any given decoder weight has minimal effect, since only their relative values are responsible for the label/output region re-mapping.